# Multimodal Clinical Trial Outcome Prediction with Large Language Models

### v3

@@ Abstract.
The clinical trial is a pivotal and costly process, often spanning multiple years and requiring substantial financial resources. Therefore, the development of clinical trial outcome prediction models aims to exclude drugs likely to fail and holds the potential for significant cost savings. Recent data-driven attempts leverage deep learning methods to integrate multimodal data for predicting clinical trial outcomes. However, these approaches rely on manually designed modal-specific encoders, which limits both the extensibility to adapt new modalities and the ability to discern similar information patterns across different modalities. To address these issues, we propose a multimodal mixture-of-experts (LIFTED) approach for clinical trial outcome prediction. Specifically, LIFTED unifies different modality data by transforming them into natural language descriptions with large language models. Then, LIFTED constructs unified noise-resilient encoders to extract information from modal-specific language descriptions. Subsequently, a sparse Mixture-of-Experts framework is employed to further refine the representations, enabling LIFTED to identify similar information patterns across different modalities and extract more consistent representations from those patterns using the same expert model. Finally, a mixture-of-experts module is further employed to dynamically integrate different modality representations for prediction, which gives LIFTED the ability to automatically weigh different modalities and pay more attention to critical information. The experiments demonstrate that LIFTED significantly enhances performance in predicting clinical trial outcomes across all three phases compared to the best baseline, showcasing the effectiveness of our proposed key components. Our code is released at https://github.com/shenmishajing/lifted.
Clinical Trial Outcome Prediction, Mixture-of-Experts, Multimodal Learning, Large Language Models
1. Introduction
@@
Figure 1. An overview of LIFTED. Step 1: Transforming multimodal data into natural language descriptions, where all modalities are converted into natural language descriptions to facilitate the representation extraction process of the transformer encoders. Step 2: Extract and combine representations from different modalities, where representations are extracted by the noise-resilient unified encoders and integrated by a Mixture-of-Experts (MoE) framework to make the final predictions.
@@
The clinical trial is a crucial step in the development of new treatments to demonstrate the safety and efficacy of the drug. Drugs must pass three trial phases involving human participants with target diseases before approval for manufacturing. However, the clinical trial is time-consuming and experiments expensive, taking multiple years and costing up to hundreds of millions of dollars (Martin et al., 2017). In addition, the success rate of clinical trials is exceedingly low and many drugs fail to pass these clinical trials (Blass, 2015; Huang et al., 2020a). Therefore, the ability to predict clinical trial outcomes beforehand, allowing the exclusion of drugs with a high likelihood of failure, holds the potential to yield significant cost savings. Given the increasing accumulation of clinical trial data over the past decade (e.g., drug descriptions, and patient criteria), we can now leverage this wealth of data for the prediction of clinical trial outcomes.
@@
Early attempts aim to improve the clinical trial outcome prediction results by modeling the components of the drugs (e.g., drug toxicity (Gayvert et al., 2016), modeled the pharmacokinetics (Qi and Tang, 2019)). Recently, deep learning methods have been proposed for trial outcome predictions. For instance, Lo et al. (2019) predicted drug approvals for 15 different disease groups by incorporating drug and clinical trial features into machine learning models. Fu et al. (2022) proposed an interaction network leveraging multimodal data (e.g., molecule information, trial documents) to capture correlations for trial outcome predictions. However, this approach relies on modal-specific encoders to extract representations from different modal data, which require manually designed encoder structures and limit their extensibility when new modal data becomes available for use.
@@
To address this, we aim to design a unified encoder to extract representations from various modalities, but it poses the following three challenges:
@@
  • How to extract representations from different modalities with a unified encoder? Different modalities are represented in various data formats. For instance, molecule information is typically depicted as a graph, while disease names rely on relationships between different diseases. Therefore, a unified encoder structure should be capable of unifying these different formats to effectively extract information.
@@
  • How to integrate extracted information from different modalities? Extracted representations from various modalities need to be integrated for predictions. However, the contribution of extracted information from different modalities may vary significantly between samples. For instance, in one patient, a specific disease, such as type 2 diabetes mellitus, which is difficult to treat, may strongly influence the final outcome. In contrast, another patient’s trial result may be primarily determined by the drugs they are prescribed, particularly if those medications have a high success rate in treating the disease. Hence, an approach to automatically weighting representations from different modalities is crucial.
@@
To address those challenges, we propose an approach called muLti-modal mIx-of-experts For ouTcome prEDiction (LIFTED), which extracts information from different modalities with a transformer based unified encoder, enhances the extracted features by a Sparse Mixture-of-Experts (SMoE) framework and integrates multimodal information with Mixture-of-Experts (MoE). Specifically, LIFTED unifies diverse multimodal features, even those in different formats, by converting them into natural language descriptions. Subsequently, we build a unified transformer-based encoder to extract representations from these modal-specific language descriptions and refine the representations with an SMoE framework. Here, the representations from different modalities are dynamically routed by a noisy top-k gating network to a portion of shared expert models, facilitating the extraction of similar information patterns. In addition, we introduce representation augmentation to enhance the resilience of transform-based encoders and the SMoE framework to potential data noise introduced during the data collection process. Furthermore, LIFTED treats the extracted representations from various modalities as distinct experts and utilizes a Mixture-of-Experts module to dynamically combine these multimodal representations for each example. This dynamic combination allows for the automatic assignment of higher weights to more crucial modalities.
Our primary contribution in this paper is LIFTED, a novel method for clinical outcome prediction. This approach unifies multimodal information through natural language descriptions and integrates multimodal information using a mixture-of-experts framework. Empirically, LIFTED is evaluated on the HINT benchamark (Fu et al., 2022) for clinical trial outcome prediction. The results demonstrate that LIFTED outperforms existing methods. Additionally, quantitative ablation studies further substantiate the effectiveness of the proposed components within LIFTED.
2. Multimodal Mixture-of-Experts for Outcome Prediction
2.1. Overview
@@
This section presents our proposed muLti-modal mIx-of-experts For ouTcome prEDiction (LIFTED) method. The goal of LIFTED is to unify multimodal data using natural language descriptions and integrate this information within a Mixture-of-Experts (MoE) framework, as illustrated in Figure 1. To elaborate, we start by extracting specific modalities from the clinical trial dataset, subsequently transforming this multimodal data into natural language descriptions using a Large Language Model (LLM). Following this, we augment the embeddings of the language descriptions derived from these different modalities. We then feed both the original and augmented embeddings into transformer-based encoders for representation learning. Subsequently, an SMoE framework is utilized to route the embeddings from different modalities to different sets of experts, where similar information patterns in different modalities will be routed to the same experts while the different patterns will be routed to experts with more specialized knowledge. To enhance the robustness of encoders, we introduce a consistency loss that aligns the original representations with the augmented ones. Moving forward, we implement an MoE framework to integrate these representations for each trial, which originate from various modalities. Finally, these integrated representations are input into a classifier for prediction. Simultaneously, we introduce an auxiliary unimodal prediction loss to improve the quality of modal-specific representations. Below, we detail LIFTED.
2.2. Transforming Multimodal Data into Natural Language Descriptions
@@
To build a unified encoder, the key challenge is how to unify multimodal data, which often have different structures for different modalities. For instance, molecule information is typically depicted as a graph, while disease names rely on relationships between different diseases. In LIFTED, we unify these different modality data by converting them into natural language descriptions. Specifically, we first format the input features into a key-value pair. After that, we use a prompt coupled with the corresponding key-value pair to ask an LLM to generate a natural language description for our input. Subsequently, these descriptions will be fed into a unified tokenizer for further encoding, except the SMILES string modality, which is tokenized by a specifically designed tokenizer to enhance the representation of molecule information. The first two steps, linearization and prompting, are detailed below:
@@
Linearization. In linearization, we format each data point $x_{i,k}$ of trial $i$ and modality $k$ into a key-value pair. In this pair, the key of each element represents the feature name $c_{i,k}$, and the corresponding value is $x_{i,k}$. This can be formulated as follows:
(1) $\mathrm{Linearize}(x_{i,k})=\{c_{i,k}:x_{i,k}\}.$
Figure 2. Processes of the linearization and the prompting. The left dashed arrow represents the linearization and the right solid arrow represents the prompting.
@@
Prompting. As depicted in Figure 2, the prompts we use to communicate with the LLM consist of three components: a prefix $p$ to describe the schema of the input features, the linearization and a suffix $s$ to instruct the LLM on how to describe the input data point in natural language. Given the prompts, the LLM will generate a readable and concise natural description $z_{i,k}$, which can be formulated as:
(2) $z_{i,k}=\mathrm{LLM}(p,\mathrm{Linearize}(x_{i,k}),s).$
@@
(3) $z_{i,0}=\mathrm{LLM}(p,$ $\;\mathrm{Linearize}(x_{i}),s),$
@@
In this way, we can summarize all the information for each clinical trial.
2.3. Representation Learning and Refinement
After transforming multimodal data into natural language descriptions, we build $K+1$ transformer-based encoders on the top on these descriptions. Specifically, each modality description $z_{i,k}$ is tokenized into a sequence of tokens $\{z_{i,k}^{t}\}_{t=1}^{T}$ with length $T$ by a tokenizer $\mathcal{T}$ and embedded into a sequence of embeddings $\{u_{i,k}^{t}\}_{t=1}^{T}$ by a modal-specific embedding layer $\mathcal{E}_{k}$ first, and then they are added by the position embeddings $\mathrm{pos}^{t}$ and fed into the correponding modal-specific transformer encoder $\mathcal{F}_{k}$ coupled with a learnable token $[cls]_{k}$ to get encoded representation $U_{i,k}$. The encoding process can be formulated as follows:
(4) $\{z_{i,k}^{t}\}_{t=1}^{T}=$ $\mathcal{T}(z_{i,k})$
@@ 2.3. Representation Learning and Refinement
Furthermore, to equip LIFTED with the capability to dynamically identify similar information patterns across different modalities and route them to the same encoder, we employ a Sparse Mixture-of-Experts (SMoE) framework to further refine the extracted representations. As illustrated in Figure 3, the encoded representations $U_{i,k}$ from different modalities will be dynamically routed by a modality-independent noisy top-k gating network $\mathcal{G}$ to a subset of shared expert models $\{\mathcal{R}^{r}\}_{r=1}^{R}$ to facilitate the extraction of similar information patterns, following the original design of SMoE (Shazeer et al., 2017). The whole process can be formulated as follows:
(5) $\mathcal{G}(U_{i,k})=$ $\mathrm{Softmax}(\mathrm{TopK}(\mathcal{P}(U_{i,k}),k))$
@@
(6) $\tilde{U}_{i,k}=$ $\mathcal{G}(U_{i,k})\cdot\mathcal{R}(U_{i,k})=$ $\sum_{r=1}^{R}\mathcal{G}^{r}(U_{i,k})\mathcal{R}^{r}(U_{i,k}).$
Figure 3. An overview of the Sparse Mixture-of-Experts framework. The encoded representations are dynamically routed by a modality-independent gating network to a subset of shared expert models. In this case, both the disease and drug name modalities utilize the same Expert 2 to extract similar information patterns.
2.4. Representation Augmentation and Consistency Loss
However, building informative modal-specific encoders and the SMoE framework solely from these modal-specific natural language descriptions remains challenging, primarily due to potential data noise introduced during the data collection process. To make the encoders and the SMoE framework more robust to the noise in the data, we augment the embeddings $u_{i,k}^{t}$ with a minor perturbation to $v_{i,k}^{t}$ and add a consistency loss to require the encoders and the SMoE framework insensitive to small perturbation, which is detailed as the following two steps:
Representation Augmentation. To perform representation augmentation, we begin by considering each embedding vector $u^{t}_{i,k}\in\mathbb{R}^{L}$, where $L$ represents the number of elements $\{m_{l}\}_{l=1}^{L}$. We randomly select a subset of these elements from $u^{t}_{i,k}$ with a probability $p$ for perturbation, while leaving the remaining elements unchanged. For simplicity, we will omit the subscript and superscript for $m_{l}$ specific to the embedding $u^{t}_{i,k}$. Next, we proceed to sample a small value $\alpha_{l}$ from a uniform distribution $\mathrm{Uniform}(-\lambda,\lambda)$ for each selected element. Here, $\lambda$ serves as a hyperparameter that controls the magnitude of the minor perturbation. Following this, each selected element is multiplied by $\exp(\alpha_{l})$ to apply the perturbation. This process can be expressed as follows:
(7) $\hat{m}_{l}=\left\{\begin{aligned} &\exp(\alpha_{l})*m_{l},&m_{l}\ \text{is selected}\\ &m_{l},&\text{otherwise}\\ \end{aligned}\right.$
After applying the perturbation, we obtain the perturbed vector $v^{t}_{i,k}=\{\hat{m}_{l}\}_{l=1}^{L}$.
Consistency Loss. After the representation augmentation step, we obtain a sequence of perturbed embeddings $\{v_{i,k}^{t}\}_{t=1}^{T}$ derived from the original embeddings $\{u_{i,k}^{t}\}_{t=1}^{T}$. These perturbed embeddings are then input into the encoder $\mathcal{F}_{k}$ and the SMoE framework to generate the encoded representation $\tilde{V}_{i,k}$. In order to ensure the robustness of the encoded embeddings, we introduce a consistency loss $\mathcal{L}_{con}$ to control the disparity between the encoded representation of the original embeddings and the augmented embeddings. This consistency loss can be formulated as follows
(8) $\mathcal{L}_{con}=\frac{1}{N(K+1)}\sum_{k=0}^{K}$ $\sum_{i=1}^{N}\left\|\tilde{U}_{i,k}-\tilde{V}_{i,k}\right\|^{2}_{F},$

### v4

# LIFTED: Multimodal Mixture-of-Experts for Clinical Trial Outcome Prediction

@@ Abstract
The clinical trial is a pivotal and costly process, often spanning multiple years and requiring substantial financial resources. Therefore, the development of clinical trial outcome prediction models aims to exclude drugs likely to fail and holds the potential for significant cost savings. Recent data-driven attempts leverage deep learning methods to integrate multimodal data for predicting clinical trial outcomes. However, these approaches rely on manually designed modal-specific encoders, which limits both the extensibility to adapt new modalities and the ability to discern similar information patterns across different modalities. To address these issues, we propose a multimodal mixture-of-experts (LIFTED) approach for clinical trial outcome prediction. Specifically, LIFTED unifies different modality data by transforming them into natural language descriptions. Then, LIFTED constructs unified noise-resilient encoders to extract information from modal-specific language descriptions. Subsequently, a sparse Mixture-of-Experts framework is employed to further refine the representations, enabling LIFTED to identify similar information patterns across different modalities and extract more consistent representations from those patterns using the same expert model. Finally, a mixture-of-experts module is further employed to dynamically integrate different modality representations for prediction, which gives LIFTED the ability to automatically weigh different modalities and pay more attention to critical information. The experiments demonstrate that LIFTED significantly enhances performance in predicting clinical trial outcomes across all three phases compared to the best baseline, showcasing the effectiveness of our proposed key components.
1 Introduction
@@
Figure 1: An overview of LIFTED. Step 1: Transforming multimodal data into natural language descriptions, where all modalities are converted into natural language descriptions to facilitate the representation extraction process of the transformer encoders. Step 2: Extract and combine representations from different modalities, where representations are extracted by the noise-resilient unified encoders and integrated by a Mixture-of-Experts (MoE) framework to make the final predictions.
@@
The clinical trial is a crucial step in the development of new treatments to demonstrate the safety and efficacy of the drug. Drugs must pass three trial phases involving human participants with target diseases before approval for manufacturing. However, the clinical trial is time-consuming and experiments expensive, taking multiple years and costing up to hundreds of millions of dollars Martin et al. (2017). In addition, the success rate of clinical trials is exceedingly low and many drugs fail to pass these clinical trials Wu et al. (2022b); Huang et al. (2020). Therefore, the ability to predict clinical trial outcomes beforehand, allowing the exclusion of drugs with a high likelihood of failure, holds the potential to yield significant cost savings. Given the increasing accumulation of clinical trial data over the past decade (e.g., drug descriptions, and patient criteria), we can now leverage this wealth of data for the prediction of clinical trial outcomes.
@@
Early attempts aim to improve the clinical trial outcome prediction results by modeling the components of the drugs (e.g., drug toxicity Gayvert et al. (2016), modeled the pharmacokinetics Qi and Tang (2019)). Recently, deep learning methods have been proposed for trial outcome predictions. For instance, Lo et al. Lo et al. (2019) predicted drug approvals for 15 different disease groups by incorporating drug and clinical trial features into machine learning models. Fu et al. Fu et al. (2022) proposed an interaction network leveraging multimodal data (e.g., molecule information, trial documents) to capture correlations for trial outcome predictions. However, this approach relies on modal-specific encoders to extract representations from different modal data, which require manually designed encoder structures and limit their extensibility when new modal data becomes available for use.
@@
To address these issues, we aim to design a unified encoder to extract representations from various modalities, but it poses the following three challenges:
@@
  • How to extract representations from different modalities with a unified encoder? Different modalities are represented in various data formats. For instance, molecule information is typically depicted as graphs, while disease names rely on relationships between diseases. Therefore, a unified encoder structure should be capable of unifying these different formats to effectively extract information.
@@
  • How to integrate extracted information from different modalities? Extracted representations from various modalities need to be integrated for predictions. However, the contribution of extracted information from different modalities may vary significantly between samples. For instance, in one patient, a specific disease, such as type 2 diabetes mellitus, which is difficult to treat, may strongly influence the final outcomes Wu et al. (2022a). In contrast, another patient’s trial result may be primarily determined by the drugs they are prescribed, particularly if those medications have a high success rate in treating the disease. Hence, an approach to automatically weighting representations from different modalities is crucial.
@@
To address those challenges, we propose an approach called muLti-modal mIx-of-experts For ouTcome prEDiction (LIFTED), which extracts information from different modalities with a transformer based unified encoder, enhances the extracted features by a Sparse Mixture-of-Experts (SMoE) framework and integrates multimodal information with Mixture-of-Experts (MoE). Specifically, LIFTED unifies diverse multimodal features, even those in different formats, by converting them into natural language descriptions. Subsequently, we build a unified transformer-based encoder to extract representations from these modal-specific language descriptions and refine the representations with an SMoE framework. Here, the representations from different modalities are dynamically routed by a noisy top-k gating network to a portion of shared expert models, facilitating the extraction of similar information patterns. In addition, we introduce representation augmentation to enhance the resilience of transform-based encoders and the SMoE framework to potential data noise introduced during the data collection process. Furthermore, LIFTED treats the extracted representations from various modalities as distinct experts and utilizes a Mixture-of-Experts module to dynamically combine these multimodal representations for each example. This dynamic combination allows for the automatic assignment of higher weights to more crucial modalities. Finally, we evaluate LIFTED on the HINT benchmark Fu et al. (2022) and the CTOD benchmark Gao et al. (2024) to demonstrate the effectiveness of LIFTED and the effectiveness of our proposed components.
2 Multimodal Mixture-of-Experts for Clinical Trial Outcome Prediction
@@
This section presents our proposed muLti-modal mIx-of-experts For ouTcome prEDiction (LIFTED) method. The goal of LIFTED is to unify multimodal data using natural language descriptions and integrate this information within a Mixture-of-Experts (MoE) framework, as illustrated in Figure 1. To elaborate, we start by extracting specific modalities from the clinical trial dataset, subsequently transforming this multimodal data into natural language descriptions using a Large Language Model (LLM). Following this, we augment the embeddings of the language descriptions derived from these different modalities. We then feed both the original and augmented embeddings into transformer-based encoders for representation learning. Subsequently, an SMoE framework is utilized to route the embeddings from different modalities to different sets of experts, where similar information patterns in different modalities will be routed to the same experts while the different patterns will be routed to experts with more specialized knowledge. To enhance the robustness of encoders, we introduce a consistency loss that aligns the original representations with the augmented ones. Moving forward, we implement an MoE framework to integrate these representations for each trial, which originate from various modalities. Finally, these integrated representations are input into a classifier for prediction. Simultaneously, we introduce an auxiliary unimodal prediction loss to improve the quality of modal-specific representations.
2.1 Transforming Multimodal Data into Natural Language Descriptions
@@
To build a unified encoder, the key challenge is how to unify multimodal data, which often have different structures for different modalities. For instance, molecule information is typically depicted as a graph, while disease names rely on relationships between different diseases Wu et al. (2022a). In LIFTED, we unify these different modality data by converting them into natural language descriptions. Specifically, we first format the input features into a key-value pair. After that, we use a prompt coupled with the corresponding key-value pair to ask an LLM to generate a natural language description for our input. Subsequently, these descriptions will be fed into a unified tokenizer for further encoding, except the SMILES string modality, which is tokenized by a specifically designed tokenizer to enhance the representation of molecule information. The first two steps, linearization and prompting, are detailed below:
Linearization.
@@
In linearization, we format each data point $x_{i,k}$ of trial $i$ and modality $k$ into a key-value pair. In this pair, the key of each element represents the feature name $c_{i,k}$, and the corresponding value is $x_{i,k}$. This can be formulated as follows:
$\mathrm{Linearize}(x_{i,k})=\{c_{i,k}:x_{i,k}\}.$ (1)
Figure 2: Processes of the linearization and the prompting.
Prompting.
@@
As depicted in Figure 2, the prompts we use to communicate with the LLM consist of three components: a prefix $p$ to describe the schema of the input features, the linearization and a suffix $s$ to instruct the LLM on how to describe the input data point in natural language. Given the prompts, the LLM will generate a readable and concise natural description $z_{i,k}$, which can be formulated as:
$z_{i,k}=\mathrm{LLM}(p,\mathrm{Linearize}(x_{i,k}),s).$ (2)
@@
$z_{i,0}=\mathrm{LLM}(p,$ $\;\mathrm{Linearize}(x_{i}),s),$ (3)
@@
By this, all the information is summarized in text format.
2.2 Representation Learning and Refinement
After transforming multimodal data into natural language descriptions, we build $K+1$ transformer-based encoders on top of these descriptions. Specifically, each modality description $z_{i,k}$ is tokenized into a sequence of tokens $\{z_{i,k}^{t}\}_{t=1}^{T}$ with length $T$ by a tokenizer $\mathcal{T}$ and embedded into a sequence of embeddings $\{u_{i,k}^{t}\}_{t=1}^{T}$ by a modal-specific embedding layer $\mathcal{E}_{k}$ first, and then they are added by the position embeddings $\mathrm{pos}^{t}$ and fed into the correponding modal-specific transformer encoder $\mathcal{F}_{k}$ coupled with a learnable token $[cls]_{k}$ to get encoded representation $U_{i,k}$. The encoding process can be formulated as follows:
$\{z_{i,k}^{t}\}_{t=1}^{T}=\mathcal{T}(z_{i,k})$ (4)
@@ 2.2 Representation Learning and Refinement
Furthermore, to equip LIFTED with the capability to dynamically identify similar information patterns across different modalities and route them to the same encoder, we employ a Sparse Mixture-of-Experts (SMoE) framework to further refine the extracted representations. The encoded representations $U_{i,k}$ from different modalities will be dynamically routed by a modality-independent noisy top-k gating network $\mathcal{G}$ to a subset of shared expert models $\{\mathcal{R}^{r}\}_{r=1}^{R}$ to facilitate the extraction of similar information patterns, following the original design of SMoE Shazeer et al. (2017). The whole process can be formulated as follows:
$\mathcal{G}(U_{i,k})=$ $\mathrm{Softmax}(\mathrm{TopK}(\mathcal{P}(U_{i,k}),k))$ (5)
@@
$\tilde{U}_{i,k}=$ $\mathcal{G}(U_{i,k})\cdot\mathcal{R}(U_{i,k})=$ $\sum_{r=1}^{R}\mathcal{G}^{r}(U_{i,k})\mathcal{R}^{r}(U_{i,k}).$ (6)
2.3 Consistent Representation Augmentation
However, building informative modal-specific encoders and the SMoE framework solely from these modal-specific natural language descriptions remains challenging, primarily due to potential data noise introduced during the data collection process. To make the encoders and the SMoE framework more robust to the noise in the data, we augment the embeddings $u_{i,k}^{t}$ with a minor perturbation to $v_{i,k}^{t}$ and add a consistency loss to require the encoders and the SMoE framework insensitive to small perturbation.
Representation Augmentation.
To perform representation augmentation, we begin by considering each embedding vector $u^{t}_{i,k}\in\mathbb{R}^{L}$, where $L$ represents the number of elements $\{m_{l}\}_{l=1}^{L}$. We randomly select a subset of these elements from $u^{t}_{i,k}$ with a probability $p$ for perturbation, while leaving the remaining elements unchanged. For simplicity, we will omit the subscript and superscript for $m_{l}$ specific to the embedding $u^{t}_{i,k}$. Next, we proceed to sample a small value $\alpha_{l}$ from a uniform distribution $\mathrm{Uniform}(-\lambda,\lambda)$ for each selected element. Here, $\lambda$ serves as a hyperparameter that controls the magnitude of the minor perturbation. Following this, each selected element is multiplied by $\exp(\alpha_{l})$ to apply the perturbation. This process can be expressed as:
$\hat{m}_{l}=\left\{\begin{aligned} &\exp(\alpha_{l})*m_{l},&m_{l}\ \text{is selected}\\ &m_{l},&\text{otherwise}\\ \end{aligned}\right.$ (7)
After his, we get the perturbed vector $v^{t}_{i,k}=\{\hat{m}_{l}\}_{l=1}^{L}$.
Consistency Loss.
After the representation augmentation step, we obtain a sequence of perturbed embeddings $\{v_{i,k}^{t}\}_{t=1}^{T}$ derived from the original embeddings $\{u_{i,k}^{t}\}_{t=1}^{T}$. These perturbed embeddings are then input into the encoder $\mathcal{F}_{k}$ and the SMoE framework to generate the encoded representation $\tilde{V}_{i,k}$. In order to ensure the robustness of the encoded embeddings, we introduce a consistency loss $\mathcal{L}_{con}$ to control the disparity between the encoded representation of the original embeddings and the augmented embeddings. This consistency loss can be formulated as follows
$\mathcal{L}_{con}=\frac{1}{N(K+1)}\sum_{k=0}^{K}$ $\sum_{i=1}^{N}\left\|\tilde{U}_{i,k}-\tilde{V}_{i,k}\right\|^{2}_{F},$ (8)